OpenStudy (lukecrayonz):

How many kilograms are in 4 moles of Na2CO3?

OpenStudy (anonymous):

Whats the molar mass of that? I don't have a periodic table so I keep asking

OpenStudy (lukecrayonz):

zbay, http://www.webqc.org/mmcalc.php for future use:)

OpenStudy (lukecrayonz):

Molar mass of Na2CO3 is 105.9886 g/mol

OpenStudy (anonymous):

\[4 mole Na_2CO_3(\frac{105.99 g Na_2CO_3}{1 mole Na_2CO_3})(\frac{1 kg}{1000 g })=.424 kg Na_2CO_3\]

OpenStudy (anonymous):

Is the equation loading or does it look all crazy for you?

OpenStudy (lukecrayonz):

Looks fine:)

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

let me first tell u how to calculate molar mass ok? u need to find of Na2CO3 so add the molar masses of the elements together multiplying them with their respective subscripts so here it goes like this Na= 23 C=12 O=16 so Na2CO3=2(23)+12+3(16)=46+12+48=96

OpenStudy (xishem):

\[46+12+48=106\]

OpenStudy (anonymous):

now 1 mole of sodium carbonate has 96 grams so 4 moles will have 4x106=424 oh yeah sry

OpenStudy (anonymous):

oops sry now that means one mole of Na2CO3 will have 106 grams

OpenStudy (anonymous):

thanks

OpenStudy (xishem):

To get the final answer, you need to convert to kg (:.

OpenStudy (anonymous):

well that is easy enough... from here the question can b done by the asker himself yhe main question has been solved isnt it? even so the answer is .106 kg ok?

OpenStudy (xishem):

Well, zbay has already answered the question. It's .424kg of Na_2CO_3

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

hah i again forgot to multiply by 4 wat a shame yeah k the answer is .424 kg i m really sry

OpenStudy (xishem):

Don't be sorry! We're all learning here.